[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun Jul 23, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun Jul 23 20:09:02 UTC 2017


HSU Shift Log for Sun Jul 23 2017 16:08:56 GMT-0400 (EDT)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues: The mirror screen of AWIPS 2 over the Atlantic  
workstation Tilefish (LXD-NHCN) is very pixelated and not clear. The screen  
on the native AWIPS2 workstation itself seems fine. Not sure what the issue  
is.

Other Techincal issues: SSMI data only showing up at FNMOC, not NRL,  
despite ATCF getting IR/vis imagery. Will investigate with Josh Cossuth.  
Just seems to be a lot of missing NRL microwave data overall.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170723/6b00c704/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list